OpenStudy (anonymous):

there is one problem i can not solve at all it is sin^2/1+cos

OpenStudy (eyust707):

simplify?

OpenStudy (anonymous):

yes i need simplify it

OpenStudy (eyust707):

do oyu know any trig identites where you get a (1 + cosx) or maybe one for sin^2x?

OpenStudy (inkyvoyd):

Think. sin(x+y)=sinxcosy+sinycosx sin(x+x)=? ??? Profit.

OpenStudy (anonymous):

the closest one is 1-sin^\[1-\sin ^2\theta\]

OpenStudy (anonymous):

@eyust707 the one i think you are suggesting is 1-sin^2 and that equals cos^2

OpenStudy (eyust707):

yes.... there is a version of that for sin^2x...

OpenStudy (lgbasallote):

i have this crazy idea....you know that \[\Huge \frac{\frac{a}{c}}{\frac{b}{c}} = \frac{a}{b}\] right? so dived the numerator and denominator by sin^2 x :DDDD

OpenStudy (anonymous):

@lgbasallote you really confused me

OpenStudy (lgbasallote):

okay..disregard what i said :P lol it's just a trick haha..i'll try to find something else

OpenStudy (inkyvoyd):

Remember, 1=sin^2+cos^2... Why don't you try putting that into the 1?

OpenStudy (anonymous):

1-cos^2/1+cos is as far as i got but i think if i was to take cos out of the top i will be left with 1-cos/1+cos then idk

OpenStudy (anonymous):

in terms of what are you supposed to simply it in?

OpenStudy (lgbasallote):

lol... sin^2 x = 1 - cos ^2x \[\huge \frac {1- cos^{2}x}{1-cos}\] \[\huge \frac{(1+cos)(1-cos)}{1-cos}\] why didn't we see it before :P

OpenStudy (anonymous):

im just suppose to get it to the simpliest form

OpenStudy (anonymous):

lol thank you its 1+ cos but that isnt the answer ugh

OpenStudy (lgbasallote):

lol then cancel the 1 + cos haha 1- cos will be left..

OpenStudy (anonymous):

im doing a puzzle with pieces and that isnt one of the answer

OpenStudy (lgbasallote):

1- cos is not the answer??

OpenStudy (anonymous):

the problems are printed on the pieces and we have to solve so im guessing that not the answer

OpenStudy (anonymous):

his answer looks correct to me, can you list what options you have?

OpenStudy (anonymous):

im trying to see if it can be simplifed again but i might have to email my teacher if it isnt too late

OpenStudy (lgbasallote):

well do you have the answer? maybe i can see what they did from the answer...what i did was.. \[\huge \frac {sin^{2}x}{1+cosx}\] \[\huge \frac{1-cos^2{x}}{1+cosx}\] \[\huge \frac{(1+cosx)(1-cosx)}{1+cosx}\] cancel 1+ cosx... 1-cosx should be the answer...i don't see how it's not :/

OpenStudy (anonymous):

that is what i got but my teacher said it should be able to simplify more bt i dnt thinkso

OpenStudy (eyust707):

nope 1-cosx is as simple as it gets

OpenStudy (anonymous):

you should get \[1-\cos(x)\] and that is it

OpenStudy (anonymous):

ok

OpenStudy (lgbasallote):

hmmm it can be simplified further actually... \[\huge 1 - cosx\] \[\huge sin^2{x} + cos^{2}x - cosx\] \[\huge sin^2{x} + cosx(cosx - 1)\] \[\large 1 - cos^2{x} + cosx(cosx-1)\] \[\large 1 - cosx(cosx + cosx - 1)\] \[\large 1 - cosx(2cosx - 1)\] lol i don't think this is simple :P

OpenStudy (anonymous):

that seems like entirely too much

OpenStudy (anonymous):

so im going to stick with 1-cosA